solve: 3 sin x - 2 = 7 sin x -1
Answer by RAY100(1637)

3 sinx -2 = 7sinx -1
.
-1 = 4 sinx
.
-1/4 =-.250 = sinx
.
taking arc sign
.
x = - 14.5 deg or 194.5 deg (180 +14.5)
.
checking
.
3(-.25) -2 = 7(-.25) -1
-2.75 = -2.75 ok